mdobbs6856:

PR = 9x-31 and QR = 43; Find x

mdobbs6856:

@smokeybrown @dude

mdobbs6856:

@pixel

SmokeyBrown:

Hi, what is the relationship between PR and QR? It seems like they are sides of some geometric shape, which might help us determine how the two are related. I don't suppose there is a diagram to go along with this question? Seeing that would certainly help to find the answer

mdobbs6856:

No diagram. I had to draw a line segment and both QR and PR are Congruent soooo...?

OrAE:

If PR (9)x-31 is your y axis the QR=43 is you x axis, But if u switch the x and y axis u would get PR as ur x and QR as your y

mdobbs6856:

That doesn't seem right. I've had to work out all these problems. so, that wouldn't make sense. @OrAE

SmokeyBrown:

@mdobbs6856 wrote:
No diagram. I had to draw a line segment and both QR and PR are Congruent soooo...?
If PR and QR are "congruent", then they will have the same length. This allows us to define the relationship: 9x - 31 = 43 First, let us add 31 to both sides of this equation: 9x = 74 Then, we can divide both sides by 9 to solve for x: x = 74/9 And there you have it, we've solved for x, using the geometric relation and basic algebra!

OrAE:

@mdobbs6856 wrote:
That doesn't seem right. I've had to work out all these problems. so, that wouldn't make sense. @OrAE
if u really take the time to think about it then it will start to make sense, trust me i had the same function yesterday in my math class

mdobbs6856:

1 attachment
mdobbs6856:

see? I put in 13 and at the top right it's green which means it's right, but idk how.

OrAE:

Well sorry 2 tell u but its wrong, like if they think thats da right answer den ok thats them but thats wrong

SmokeyBrown:

@mdobbs6856 wrote:
see? I put in 13 and at the top right it's green which means it's right, but idk how.
That's interesting. I've used programs before where you get a green light or a checkmark just for putting *anything* in the answer space, whether it's correct or not. Are you sure that is being accepted as the *correct* answer, or merely *an* answer? You could do a quick experiment: try to fill the space with an answer that you *know* is wrong; see if the light stays green. That will tell you whether or not a green light signifies a correct answer. I have my doubts. It does not seem like the math checks out...

mdobbs6856:

I have done that, but I'll show you

mdobbs6856:

1 attachment
SmokeyBrown:

@mdobbs6856 wrote:
I have done that, but I'll show you
Thank you for indulging me. Well, if that is the answer that is recognized as "correct", then I see nothing wrong with using that for your submission. Out of curiosity, is 74/9 accepted as a correct answer for that one? Perhaps there is some sort of typo or other error, or perhaps the error is on my part?

mdobbs6856:

1 attachment
mdobbs6856:

Thanks @smokeybrown and @OrAE

SmokeyBrown:

Aye, no worries, it looks like you did most of the work yourself, to be honest!

OrAE:

Anytime @mdobbs6856